# Enhancing LLM-based Code Translation in Repository Context via Triple Knowledge-Augmented

### v2

@@ Abstract.
Large language models (LLMs) have behaved well in function-level code translation without repository-level context. However, the performance of LLMs in repository-level context code translation remains suboptimal due to complex dependencies and context, hindering their adoption in industrial settings. In this work, we propose a novel LLM-based code translation technique K3Trans, which leverages triple knowledge augmentation to enhance LLM’s translation quality under repository context in real-world software development. First, K3Trans constructs a translation knowledge base by extracting relevant information from target-language codebases, the repository being translated, and prior translation results. Second, for each function to be translated, K3Trans retrieves relevant triple knowledge, including target-language code samples, dependency usage examples, and successful translation function pairs, serving as references to enhance LLM for translation. Third, K3Trans constructs a knowledge-augmented translation prompt using the retrieved triple knowledge and employs LLMs to generate the translated code while preserving repository context. It further leverages LLMs for self-debugging, enhancing translation correctness.
The experiments show that K3Trans substantially outperforms the baseline adapted from previous work by 19.4%/40.2% relative improvement in pass@1 and 0.138 in CodeBLEU. Furthermore, the code generated by K3Trans is of higher quality, as indicated by the higher DSR@1 metric of 82.1% and Repairable Ratio of 44.6%, which suggests a greater proportion of fixable code. It is important to note that the results also demonstrate that each knowledge significantly contributes to K3Trans ’s effectiveness in handling repository-level context code translation, with dependency usage examples making the most notable contribution. Moreover, as the self-evolution process progresses, the knowledge base continuously enhances the LLM’s performance across various aspects of the repository-level code translation.
copyright: acmlicensedjournalyear: 2024
@@ 1. Introduction
Figure 1. Motivation Examples
@@
Code translation plays a critical role in modern software development, as manual translation is both time-consuming and error-prone, necessitating robust automated tools (Jiao et al., 2023). Recent advances in large language models (LLMs) have demonstrated remarkable performance in various natural language processing tasks, including code translation (Pan et al., 2024; Yang et al., 2024; Tao et al., 2024). Conventional benchmarks, such as CodeTransOcean (Yan et al., 2023), have shown that LLMs perform well on function-level translation tasks. However, these benchmarks largely ignore the challenges posed by repository-level context, where real-world code translation must consider complex dependencies, project-specific coding styles, and evolving repository structures (Ou et al., 2024; Du et al., 2023; Zhang et al., 2025a; Wang et al., 2024). As a result, the performance of LLMs in repository-context translation (e.g., RustRepoTrans (Ou et al., 2024)) remains below expectations, limiting their practical adoption in industrial applications. Ou et al. (Ou et al., 2024) demonstrates that even the best model experiences a 30.8% performance drop (Pass@1 from 74.3% to 43.5%) when handling repository-level context compared to previous benchmarks without such context.
@@
In real-world software development, code translation is not performed on isolated functions. Instead, developers translate code incrementally, maintaining the project’s architecture while managing intricate interdependencies. Figure 1 illustrates three translation examples using Basic LLM from the RustRepoTrans benchmark that highlight common translation failures. (1) Dependency Misuse: An LLM fails to correctly apply dependency usage, leading to errors. Injecting a relevant dependency usage example from the repository into the translation prompt resolves this issue. (2) Syntax Misinterpretation: LLM produces code that violates target language syntax (e.g., error “cannot borrow ‘p’ as mutable, as it is not declared as mutable”). Providing an example from an open-source project with similar syntax characteristics guides the LLM toward correct syntax. (3) Syntactical Difference Confusion: An LLM fails to recognize syntactical differences between the source language and the target language, resulting in unnecessary checks, such as redundant null pointer or memory safety validations, which are necessary in the source language, but do not align with the target language’s paradigms. A reference to successful translation function pairs, where such checks are omitted, enables the LLM to avoid these pitfalls. Existing work (Bhattarai et al., 2024) using RAG to enhance LLM code translation has only provided code translation pairs as a single type of knowledge. It doesn’t take into account repository context information in real-world software development scenarios and the fact that in real translation tasks, the target language is often an emerging, low-resource language (Ou et al., 2024). As a result, previous approaches overlook the importance of dependency usage knowledge and target language syntax knowledge, resulting in poor performance in real translation scenarios.
@@
Motivated by these challenges, this work proposes K3Trans, a novel LLM-based code translation technique that leverages triple knowledge augmentation to enhance translation quality under repository context. Specifically, K3Trans integrates: (1) Target-language Code Samples drawn from existing projects, which provide real-world syntax and style references; (2) Dependency Usage Examples extracted from the repository being translated, offering insights into proper dependency invocation. (3) Successful Translation Function Pairs from translation history, serving as concrete examples for correct code transformation.
@@
By building a continuously updating dynamic knowledge base, K3Trans adapts to evolving repository contexts and improves the LLM’s ability to generate accurate, context-aware code. Our experiments, conducted on the newly introduced RustRepoTrans benchmark, demonstrate that incorporating triple knowledge not only mitigates common translation errors but also significantly enhances overall translation performance. We comprehensively evaluate K3Trans on the repository-level context code translation benchmark RustRepoTrans. First, we compare K3Trans with LLM-based with PE and Basic RAG adapted from (Bhattarai et al., 2024), the results show that K3Trans outperforms the baselines in all both match-based metrics and execution-based metrics. Specifically, compared to Basic RAG, the Pass@1 metric surges from 48.3% to 67.7%, while the CodeBLEU score improves from 0.595 to 0.733. Furthermore, even prompt design in K3Trans alone can achieve slightly better results than the baseline adapted from previous work. Second, we explore the impact of different types of knowledge in the knowledge base on K3Trans by progressively reducing the provided knowledge categories. The results demonstrate that each knowledge significantly contributes to K3Trans ’s effectiveness in handling repository-level context code translation, with dependency usage examples making the most notable contribution. Third, we evaluate the impact of the self-evolution process on K3Trans and find that as the self-evolution process progresses, the knowledge base continuously enhances the LLM’s performance across various aspects of the repository-level code translation.
In summary, our contributions are as follows, data and code will be publicly available later.
@@
  • Extensive Evaluation: We conduct comprehensive experiments on the RustRepoTrans benchmark, demonstrating significant improvements over existing methods in accuracy and robustness.
@@ 2.1. Overview
To improve code translation within a repository context, this work introduces K3Trans, an LLM-based technique that enhances translation through triple knowledge augmentation from three key sources: (1) target-language code samples from existing projects, (2) dependency usage examples from the repository being translated, and (3) successful translation function pairs from code translation history. Figure 2 presents the framework of K3Trans. Given a repository-level context function to translate, K3Trans leverages triple knowledge to generate target-language code. The framework consists of three phases: Phase-1 (Translation Knowledge Base Construction) is performed offline, whereas Phase-2 (Translation Knowledge Retrieval) and Phase-3 (Knowledge-Augmented Code Translation) operate online.
@@
  • Phase-2: Translation Knowledge Retrieval (Section 2.3). For each function to be translated, K3Trans retrieves relevant triple knowledge, including target-language code samples, dependency usage examples, and successful translation function pairs, serving as references for translation.
@@
K3Trans operates as a self-evolving framework, continuously improving translation quality as target-language codebases, repository structures, and translation history evolve over time. By dynamically incorporating up-to-date translation knowledge, K3Trans enhances the adaptability and accuracy of code translation.
@@ 2.2. Translation Knowledge Base Construction
Figure 3. Overview of Prompt Construction (using C to Rust as example)
In K3Trans, the translation knowledge base consists of three parts: target-language code samples from projects in target language (in Section 2.2.1), dependency usage examples from target projects (in Section 2.2.2), and translation pair examples from previous successful translation (in Section 2.2.3), which are aimed to enhance the LLM’s understanding of the target language’s syntax, its ability to correctly recognize and invoke dependencies, and its capability to identify syntax differences.
@@ 2.2.1. Target-language Code Samples Extraction
Existing work has shown that providing the model with code snippets similar to the target code as a reference improves the quality of the generated code in code generation tasks (Fan et al., 2024; Du et al., 2024). Since code implementing the same functionality often uses similar function names and variable names, leading to higher textual similarity (Zhong et al., 2010), retrieving the target-language code with the highest textual similarity to the source function allows us to effectively identify code samples from similar-domain projects that implement comparable functionality. These code samples could help improve translation as well. To simulate the most comprehensive reference code for the target language currently available, we obtained open source projects on GitHub that were created before March 31, 2022, from the The Stack dataset (Kocetkov et al., 2022) to directly assess function-level code samples. Then, using the official GitHub API, we crawled open source projects created after March 31, 2022, based on the condition ‘language: \$target_language created:2022-04-01..2025-02-01’, and then used tree-sitter (tree sitter, 2023) to extract function from source file. To prevent potential data leakage in the knowledge base, we remove all projects with the same name as those involved in evaluation benchmark (Ou et al., 2024), regardless of whether their creators are the same, considering the possibility of project forks. These two parts together form the General Projects of the knowledge base, containing a total of 113,882,369 target language code examples from 671,240 projects, which are used to enhance the code generated by the LLM to better conform to the syntax requirements of the target language.
@@
2.2.2. Dependency Usage Examples Extraction
Since the same dependency is often called multiple times within the same project, the way a dependency is called elsewhere in the project can serve as a useful reference for its usage in the target function. Specifically, when the scope is the same, the same dependency will have the same call path and identical types of input parameters. When the scope differs, the call path will be similar and the input parameters will still be of the same type. Therefore, for each dependency involved in a translation task, we extract the corresponding dependency call statements from the code in the current project that shares the same scope as the target function. If multiple call statements are found, only the first occurrence is retained as the example. Specifically, by utilizing tree-sitter (tree sitter, 2023), we extract function dependencies by identifying call_expression nodes to retrieve all function invocation statements. These statements are then filtered based on the names of the target function dependency, obtaining the required dependency invocation statements. For variable dependencies, we extract code execution statements and filter them according to the names of the target variable dependency, thereby obtaining the corresponding variable dependency invocation statements. Ultimately, for each dependency, we constructed the corresponding dependency usage examples with [source code of dependency] and [usage example] (if no corresponding invocation statement was retrieved, [usage example] is set to None). These extracted statements form the dependency usage examples, enhancing the LLM’s ability to correctly identify and call dependencies.
This component is self-evolving by continuously performing the previously mentioned dependency extraction operations during the translation process.
@@ 2.2.3. Successful Translation Function Pair Examples Collection
Providing the model with information of the same category as the target task could maximize its performance on the current task. However, due to the scarcity of existing repository-level context code translation datasets (Ou et al., 2024), it is difficult to obtain a large number of high-quality function-level equivalence pairs to construct successful translation function pairs base, for repository-level context. Therefore, we collect successful translation function pairs during the translation process and add them to the knowledge base, constructing the successful translation function pairs knowledge base, which are used to enhance the model’s ability to recognize syntactic differences when performing code translation tasks.
@@ 2.3. Translation Knowledge Retrieval
The input of the repository-level context code translation task consists of three parts: the source code of the function under translation, the function signature of target function, and the target function dependencies. For a given repository-level context code translation task, K3Trans retrieves relevant translation knowledge from the constructed translation knowledge to construct translation prompt (as shown in 3) base in a three-step retrieval process: focal function and dependencies extraction, candidate knowledge retrieval, candidate knowledge re-ranking.
@@ 2.3.3. Candidate knowledge Re-ranking
Although BM25 performs excellently in text-based similarity calculations, it struggles to effectively identify noise, such as comments, and other critical information within code, such as the Abstract Syntax Tree (AST). Therefore, K3Trans employs UniXcoder (Guo et al., 2022), a unified cross-modal pre-trained model for programming languages, to re-rank the retrieved knowledge items by calculating cosine similarity. The top n, where n = 2 for target-language code sample (will be discussed further in Section 2) and n = 1 for successful translation function pairs in our experiments candidate knowledge items with the highest UniXcoder scores are selected as the final knowledge items to be provided to the LLM-based code translation.

### v3

# K3Trans: Evolving Triple Knowledge-Augmented LLMs for Code Translation in Repository Context

@@ Abstract.
Large language models (LLMs) have behaved well in function-level code translation without repository-level context. However, the performance of LLMs in repository-level context code translation remains suboptimal due to complex dependencies and context, hindering their adoption in industrial settings. In this work, we propose a novel LLM-based code translation technique K3Trans, which leverages triple knowledge augmentation to enhance LLM’s translation quality under repository context in real-world software development. First, K3Trans constructs an evolving translation knowledge base by extracting relevant information from target-language codebases, the repository being translated, and prior translation results. Second, for each function to be translated, K3Trans retrieves relevant triple knowledge, including target-language code samples, dependency usage examples, and successful translation function pairs, serving as references to enhance LLM for translation. Third, K3Trans constructs a knowledge-augmented translation prompt using the retrieved triple knowledge and employs LLMs to generate the translated code while preserving repository context. It further leverages LLMs for self-debugging, enhancing translation correctness. Lastly, K3Trans continuously evolves the translation knowledge base by extracting new dependency usage examples and successful translation function pairs from prior translation history or from components that have been refactored, or updated manually by developers and extracting target language code samples by automatically detecting and downloading newly added target-language project.
The experiments show that K3Trans substantially outperforms the baseline adapted from previous work, achieving relative improvements of up to 135.9% on Pass@1 and 32.8% on CodeBLEU among studied LLMs. Furthermore, the code generated by K3Trans is of higher quality, as indicated by the higher DSR@1 and Repairable Ratio, which suggests a greater proportion of fixable code. It should be noted that the results also demonstrate that each knowledge significantly contributes to K3Trans ’s effectiveness in handling repository-level context code translation, with dependency usage examples making the most notable contribution. Moreover, as the self-evolution process progresses, the knowledge base continuously enhances the LLM’s performance across various aspects of the repository-level code translation.
Repository-level Context, Incremental Code Translation, Large Language Models
$c$s: Software and its engineering Model-driven software engineering
@@ 1. Introduction
Figure 1. Motivational Examples of LLMs’ Performance With vs. Without Knowledge in Addressing Three Key Challenges: Dependency Misuse, Syntax Misinterpretation, and Syntactical Difference Confusion
@@
Code translation plays a critical role in modern software development, as manual translation is both time-consuming and error-prone, necessitating robust automated tools (Jiao et al., 2023). Recent advances in large language models (LLMs) have demonstrated remarkable performance in various natural language processing tasks, including code translation (Pan et al., 2024; Yang et al., 2024; Tao et al., 2024). Conventional benchmarks, such as CodeTransOcean (Yan et al., 2023), have shown that LLMs perform well on function-level translation tasks.
However, these benchmarks largely ignore the challenges posed by repository-level context, where real-world code translation must consider complex dependencies, project-specific coding styles, and evolving repository structures (Ou et al., 2024; Zhang et al., 2025; Wang et al., 2024). Furthermore, in real-world projects, translation rarely rewrites an entire repository at once. Developers typically incrementally translate new or modified modules from the source to the target language (Zhang et al., 2023; Copilot, 2024; Mateus et al., 2023; Brant and Roberts, 2009). Newly translated code must integrate with existing modules, including components migrated earlier or manually refined, to preserve functional correctness and consistency. Incremental workflows also arise when the source project evolves after partial migration, requiring new features to be ported into the target repository. Such repository-level migration demands careful handling of non one-to-one mapping repository-level context such as different dependencies, distinct function signatures, and architectural divergence. As a result, the performance of LLMs in incremental translation scenario with evolving repository-context remains below expectations, limiting their practical adoption in industrial applications. Ou et al. (Ou et al., 2024) demonstrates that even the best model experiences a 30.8% performance drop (Pass@1 from 74.3% to 43.5%) when handling repository-level context compared to previous benchmarks without such context.
@@
Figure 1 illustrates three translation examples drawn from the evaluation results of LLMs on the RustRepoTrans benchmark that highlight common translation challenges. (1) Dependency Misuse: An LLM fails to correctly apply dependency usage, leading to errors. Injecting a relevant dependency usage example from the repository into the translation prompt resolves this issue. (2) Syntax Misinterpretation: LLM produces code that violates target language syntax (e.g., error “cannot borrow ‘p’ as mutable, as it is not declared as mutable”). Providing an example from an open-source project with similar syntax characteristics guides the LLM toward correct syntax. (3) Syntactical Difference Confusion: An LLM fails to recognize syntactical differences between the source language and the target language, resulting in unnecessary checks, such as redundant null pointer or memory safety validations, which are necessary in the source language, but do not align with the target language’s paradigms. A reference to successful translation function pairs, where such checks are omitted, enables the LLM to avoid these pitfalls. Existing work (Bhattarai et al., 2024) using RAG to enhance LLM code translation has only provided code translation pairs as a single type of knowledge. It does not consider repository context information in real-world software development scenarios or the fact that the target language in practical translation tasks is often an emerging, low-resource language (Ou et al., 2024). Furthermore, the retrieval scope of existing methods is limited to a static knowledge base, which does not evolve dynamically as the translation process advances. Consequently, previous approaches overlook the importance of dependency usage knowledge and target language syntax knowledge, resulting in poor performance in real translation scenarios. Additionally, the static nature of the knowledge base restricts the range of tasks that RAG can support, making it ineffective in handling continuously evolving repository-level context.
@@
Motivated by these challenges, this work proposes K3Trans, a novel LLM-based code translation technique that leverages evolving triple knowledge augmentation to enhance translation quality under repository context. Specifically, K3Trans integrates: (1) Dependency Usage Examples extracted from the repository being translated, offering insights into proper dependency invocation; (2) Target-language Code Samples drawn from existing projects, which provide real-world syntax and style references; (3) Successful Translation Function Pairs from translation history, serving as concrete examples for correct code transformation.
@@
By building a continuously updating dynamic knowledge base, K3Trans adapts to evolving repository contexts and improves the LLM’s ability to generate accurate, context-aware code. Our experiments, conducted on the newly repository-level context code translation benchmark RustRepoTrans, demonstrate that incorporating triple knowledge not only mitigates common translation errors but also significantly enhances overall translation performance. First, we compare K3Trans with LLM-based baselines: PE and Basic RAG adapted from (Bhattarai et al., 2024), and the Rule-based baseline: C2Rust (immunant, 2025). The results show that K3Trans outperforms all the baselines across all match-based and execution-based metrics. Specifically, compared to Basic RAG, K3Trans achieves relative improvements of up to 135.9% in Pass@1 and 32.8% in CodeBLEU among the studied LLMs. Furthermore, even the prompt design in K3Trans alone slightly outperforms the baseline adapted from prior work. Moreover, C2Rust only successfully translated a single case out of 145 tasks, demonstrating that the rule-based baseline is incapable of handling non one-to-one mapping repository-level context, which is common in real-world scenarios. Second, we explore the impact of different types of knowledge in the knowledge base on K3Trans by progressively reducing the provided knowledge categories. The results demonstrate that each knowledge significantly contributes to K3Trans ’s effectiveness in handling repository-level context code translation, with dependency usage examples making the most notable contribution. Third, we evaluate the impact of the self-evolution process on K3Trans and find that as the self-evolution process progresses, the knowledge base continuously enhances the LLM’s performance across various aspects of the repository-level code translation. Lastly, we manually analyze the successful examples of K3Trans that the baselines fail, and the failure examples of K3Trans to fully understand the advantages and limitations of K3Trans.
In summary, our contributions are as follows:
@@
  • Extensive Evaluation: We conduct comprehensive experiments on the RustRepoTrans benchmark, demonstrating significant improvements among all studied LLMs over existing methods in accuracy and robustness.
@@ 2.1. Overview
To improve code translation within a repository context, this work introduces K3Trans, an LLM-based technique that enhances translation through triple knowledge augmentation from three key sources: (1) dependency usage examples from the repository being translated, (2) target-language code samples from existing projects, and (3) successful translation function pairs from code translation history. Figure 2 presents the framework of K3Trans. Given a repository-level context function to translate, K3Trans leverages triple knowledge to generate target-language code. The framework consists of three phases: Phase-1 (Translation Knowledge Base Construction) is performed offline, whereas Phase-2 (Translation Knowledge Retrieval) and Phase-3 (Knowledge-Augmented Code Translation) operate online.
@@
  • Phase-2: Translation Knowledge Retrieval (Section 2.3). For each function to be translated, K3Trans retrieves relevant triple knowledge, including dependency usage examples, target-language code samples, and successful translation function pairs, serving as references for translation.
@@
K3Trans operates as a self-evolving framework, continuously improving translation quality as target-language codebases, repository structures, and translation history evolve over time. By dynamically incorporating up-to-date translation knowledge, K3Trans enhances the adaptability and accuracy of translation.
@@ 2.2. Translation Knowledge Base Construction
In K3Trans, the translation knowledge base consists of three parts: dependency usage examples from the repository being translated (in Section 2.2.1), target-language code samples from existing projects (in Section 2.2.2), and translation pair examples from previous successful translation (in Section 2.2.3), which are aimed to enhance LLM’s ability to correctly recognize and invoke dependencies, understanding of target language’s syntax, and capability to identify syntax differences.
2.2.1. Dependency Usage Examples Extraction
Since the same dependency is often called multiple times within the same project, the way a dependency is called elsewhere in the project can serve as a useful reference for its usage in the target function. Specifically, when the scope is the same, the same dependency will have the same call path and identical types of input parameters. When the scope differs, the call path will be similar and the input parameters will still be of the same type. Therefore, for each dependency involved in a translation task, we extract the corresponding dependency call statements from the code in the current project that shares the same scope as the target function. If multiple call statements are found, only the first occurrence is retained as the example. Specifically, by utilizing tree-sitter (tree sitter, 2023), we extract function dependencies by identifying call_expression nodes to retrieve all function invocation statements. These statements are then filtered based on the names of the target function dependency, obtaining the required dependency invocation statements. For variable dependencies, we extract code execution statements and filter them according to the names of the target variable dependency, thereby obtaining the corresponding variable dependency invocation statements. Ultimately, for each dependency, we constructed the corresponding dependency usage examples with [source code of dependency] and [usage example] (if no corresponding invocation statement was retrieved, [usage example] is set to empty). These extracted statements form the dependency usage examples, enhancing the LLM’s ability to correctly identify and invoke dependencies.
This component is self-evolving by continuously performing the dependency extraction operations mentioned before to extract dependency usage examples from successfully translated code or from components that have been refactored, or updated manually by developers that has been subsequently integrated into the project’s repository context during translation process.
@@ 2.2.2. Target-language Code Samples Extraction
Existing work has shown that providing the model with code snippets similar to the target code as a reference improves the quality of generated code in code generation tasks (Fan et al., 2024). Since code implementing the same functionality often uses similar function names and variable names, leading to higher textual similarity (Zhong et al., 2010), retrieving the target-language code with the highest textual similarity to the function under translation allows us to effectively identify code samples from similar-domain projects that implement comparable functionality. These code samples could help improve translation as well. To simulate the most comprehensive reference code for the target language currently available, we obtained open source projects on GitHub that were created before March 31, 2022, from the The Stack dataset (Kocetkov et al., 2022) to directly assess function-level code samples. Then, using the official GitHub API, we crawled open source projects created after March 31, 2022, based on the condition ‘language: \$target_language created:2022-04-01..2025-02-01’, and then used tree-sitter (tree sitter, 2023) to extract function from source file. To prevent potential data leakage in the knowledge base, we remove all projects with the same name as those involved in evaluation benchmark (Ou et al., 2024), regardless of whether their creators are the same, considering the possibility of project forks. These two parts together form the General Projects of the knowledge base, containing a total of 113,882,369 target language code examples from 671,240 projects, which are used to enhance the code generated by the LLM to better conform to the syntax requirements of the target language.
@@ 2.2.3. Successful Translation Function Pair Examples Collection
Providing the model with information of the same category as the target task could maximize its performance on the current task. However, due to the scarcity of existing repository-level context code translation datasets (Ou et al., 2024), it is difficult to obtain a large number of high-quality function-level equivalence pairs to construct successful translation function pairs base, for repository-level context. Therefore, we collect successful translation function pairs during the translation process and add them to the knowledge base, constructing the successful translation function pairs knowledge base, which are used to enhance model’s ability to recognize syntactic differences when performing code translation tasks.
@@ 2.3. Translation Knowledge Retrieval
The input of the repository-level context code translation task consists of three parts: the source code of the function under translation, the function signature of target function, and the target function dependencies. For a given repository-level context code translation task, K3Trans retrieves relevant translation knowledge from the constructed translation knowledge to construct translation prompt (as shown in Figure 3) base in a three-step retrieval process: focal function and dependencies extraction, candidate knowledge retrieval, candidate knowledge re-ranking.
Figure 3. Overview of Prompt Construction (using C to Rust as example)
@@ 2.3.3. Candidate knowledge Re-ranking
Although BM25 performs excellently in text-based similarity calculations, it struggles to effectively identify noise, such as comments, and other critical information within code, such as the Abstract Syntax Tree (AST). Therefore, K3Trans employs UniXcoder (Guo et al., 2022), a unified cross-modal pre-trained model for programming languages, to re-rank the retrieved knowledge items by calculating cosine similarity. The top n, where n = 2 for target-language code sample (will be discussed further in Section 4.3) and n = 1 for successful translation function pairs in our experiments candidate knowledge items with the highest UniXcoder scores are selected as the final knowledge items to be provided to the LLM-based code translation.